Middlefield Motors is evaluating project Z. The project would require an initial investment of 76,000 dollars that would be depreciated to 6,000 dollars over 6 years using straight-line depreciation. The first annual operating cash flow of 28,000 dollars is expected in 1 year, and annual operating cash flows of 28,000 dollars are expected each year forever. Middlefield Motors expects the project to have an after-tax terminal value of 341,000 dollars in 4 years. The tax rate is 20 percent. What is (X+Y)/Z if X is the project’s relevant expected cash flow for NPV analysis in year 4, Y is the project’s relevant expected cash flow for NPV analysis in year 5, and Z is the project’s relevant expected cash flow for NPV analysis in year 3? Round your answer to 2 decimal places (for example, 2.89, 0.70, or 1.00).
Answer #1

Since operating cash flow is already given, the tax and depreciation data is already taken into account.

Hence Y = 28000

Z= 28000

X = OCF+After tax terminal value = 28000+341000 = 369000

(X+Y)/Z = (369000+28000)/28000

=14.18